### v2

# NashOpt - A Python Library for Computing Generalized Nash Equilibria

Abstract
NashOpt is an open-source Python library for computing and designing generalized Nash equilibria (GNEs) in noncooperative games with shared constraints and real-valued decision variables. The library exploits the joint Karush-Kuhn-Tucker (KKT) conditions of all players to handle both general nonlinear GNEs and linear–quadratic games, including their variational versions. Nonlinear games are solved via nonlinear least-squares formulations, relying on JAX for automatic differentiation. Linear–quadratic GNEs are reformulated as mixed-integer linear programs, enabling efficient computation of multiple equilibria. The framework also supports inverse-game and Stackelberg game-design problems. The capabilities of NashOpt are demonstrated through several examples, including noncooperative game-theoretic control problems of linear quadratic regulation and model predictive control. The library is available at https://github.com/bemporad/nashopt.
1 Introduction
Generalized games describe situations where several agents take decisions by optimizing individual objective functions under common constraints on their decision variables. Examples include traffic participants using the same roads, wireless systems sharing limited bandwidth, energy systems where prosumers share transmission capacity and balance constraints [10, 38], or firms competing in a market [3]. The solution concept for these problems, in which the cost function and the feasible set of each agent depends on the decisions taken by the other agents, is the generalized Nash equilibrium (GNE), a state in which no agent has an incentive or the possibility to change their decision [2].
GNEPs have been widely studied in the literature, with various algorithms proposed for their solution, including splitting-based operator frameworks [39, 3], approaches built on augmented Lagrangian formulations [34, 27], interior-point optimization strategies [12], Newton-type algorithms [16], distributed algorithms [39, 37, 3], and learning-based methods only relying on measuring best responses [14]. For the specific class of GNEPs in which the shared constraints are linear, efficient solution methods exist for linear objectives [13, 20], multiparametric quadratic programming for strictly convex quadratic objectives [24], and distributed payoff-based algorithms for convex games [36].
1.1 Contribution
Motivated by the relative scarcity of software packages for solving rather general classes of GNEPs and their application to game-theoretic control problems, in this paper we describe the mathematical formulations of our package NashOpt, a Python library for solving GNEPs based on satisfying the KKT conditions of each player jointly. The library supports nonlinear GNEPs, which are solved via nonlinear least-squares methods, and linear-quadratic GNEPs, which can be solved more efficiently to global optimality via mixed-integer linear programming (MILP) and can be used to enumerate multiple equilibria for the same game corresponding to different combinations of active constraints at optimality. We also address game-design problems, where a central authority can optimize some parameters of the game to achieve a desired equilibrium in accordance with the Single-Leader-Multi-Follower (SLMF) Stackelberg game setting [1]. Moreover, we leverage the GNE solution methods to solve game-theoretic control problems, including non-cooperative linear quadratic regulation (LQR) and model predictive control (MPC) problems [23, 29]. The main purpose of the tool is to analyze and predict the effect of different agents pursuing their own objectives in a shared environment, and to design the game to promote desired equilibrium behaviors.
For nonlinear GNEs, we leverage JAX [8] for automatic differentiation and just-in-time compilation to achieve high performance, and on the open-source MILP solver HiGHS [26] or Gurobi [22] for the linear-quadratic case, using the so-called “big-M” formulation to encode the complementary slackness conditions of the KKT system [25, 19].
The library is open-source and available at https://github.com/bemporad/nashopt, and can be easily installed via pip install nashopt.
1.2 Notation
We denote by $\mathbb{R}$ the set of real numbers. Given $N$ vectors $x_{i}\in\mathbb{R}^{n_{i}}$, $i=1,\ldots,N$, we denote by $x=\mathop{\rm col}\nolimits(x_{1},\ldots,x_{N})\in\mathbb{R}^{n}$ their vertical concatenation, where $n=\sum_{i=1}^{N}n_{i}$. We denote by $\mathbf{1}$ the vector of all ones of appropriate dimension, by $I$ the identity matrix, and by $e_{i}$ the $i$th column of $I$. Given a matrix $A$, we denote by $A^{\top}$ its transpose and by $\|A\|_{F}$ its Frobenius norm.
2 Problem Formulation
Consider the following game-design problem based on a multiparametric generalized Nash equilibrium problem (GNEP) with $N$ players:
[…]
where $x_{i}\in\mathbb{R}^{n_{i}}$ collects the decision variables optimized by agent $i$, $i=1,\ldots,N$, $x_{-i}$ collects those of all other agents, $x=\mathop{\rm col}\nolimits(x_{1},\ldots,x_{N})\in\mathbb{R}^{n}$ is the stacked vector of all agents’ decision variables, $x^{\star}(p)$ is the generalized Nash equilibrium (GNE) of the game for a given parameter vector $p\in\mathbb{R}^{n_{p}}$, $n_{p}\geq 0$, where by $n_{p}=0$ we mean that no parameter enters (1). In (1), $f_{i}:\mathbb{R}^{n}\to\mathbb{R}$ is the objective of agent $i$, $g:\mathbb{R}^{n}\to\mathbb{R}^{n_{g}}$ and $h:\mathbb{R}^{n}\to\mathbb{R}^{n_{h}}$ define, respectively, (shared) inequality and equality constraints, and $\ell_{i}\in(\mathbb{R}\cup\{-\infty\})^{n_{i}}$, $u_{i}\in(\mathbb{R}\cup\{+\infty\})^{n_{i}}$ define the (local) box constraints for each player $i$. Moreover, $J:\mathbb{R}^{n}\times\mathbb{R}^{n_{p}}\to\mathbb{R}$ is the objective function used to shape the desired equilibrium behavior of the game, and $P\subseteq\mathbb{R}^{n_{p}}$ is a given feasible set for the parameter vector $p$, such as the hyper-box $P=\{p\in\mathbb{R}^{n_{p}}:\ \ell_{p}\leq p\leq u_{p}\}$ where $\ell_{p},u_{p}$ are (possibly infinite) bounds. Problem (1) is also referred to as an optimistic equilibrium of Single-Leader-Multi-Follower (Stackelberg) game [1], with the leader optimizing $p$ and the followers playing the GNEP parameterized by $p$.
A GNE $x^{\star}(p)$ is a vector for which no agent can reduce their cost given the strategies of the other players and feasibility constraints, i.e.,
$f_{i}(\mathop{\rm col}\nolimits(x_{i}^{\star},x_{-i}^{\star}),p)\leq f_{i}(\mathop{\rm col}\nolimits(x_{i},x_{-i}^{\star}),p)\qquad\forall x_{i}\in X_{i}(x_{-i}^{\star},p)$
where, with a slight abuse of notation, we denoted by $\mathop{\rm col}\nolimits(x_{i},x_{-i})$ the vector obtained by collecting the components of $x_{i}$ and $x_{-i}$ in the right order, and $X_{i}(x_{-i}^{\star},p)\subseteq\mathbb{R}^{n_{i}}$ is the feasible set of player $i$ defined by the constraints in (1) given the strategies $x_{-i}^{\star}(p)$ of the other players. We highlight special cases of (1) in the next sections.
2.1 Solving a specific GNEP
To just solve a given game, we set $J(x^{\star},p)\equiv 0$ and $P=\{p_{0}\}$ a singleton, leading to the standard GNEP:
$\begin{array}[]{rl}x_{i}^{\star}=\arg\min_{\ell_{i}\leq x_{i}\leq u_{i}}&f_{i}(x,p_{0})\\ \text{s.t. }&g(x,p_{0})\leq 0,\ h(x,p_{0})=0\\ &x_{-i}=x_{-i}^{\star},\quad i=1,\ldots,N.\end{array}$ (2)
2.2 Finding a parameter vector for which a GNE exists
If no specific equilibrium is desired, we can set $J(x^{\star},p)\equiv 0$ and let $P$ be a non-singleton set, leading to the problem of finding a parameter vector $p\in P$ for which an equilibrium exists.
2.3 Inverse game problem
Given an observed agents’ equilibrium $x_{\textrm{d}es}$ (or a desired one we wish to achieve), by setting $J(x^{\star},p)=\|x^{\star}-x_{\rm des}\|_{2}^{2}$ (or $J(x^{\star},p)=\|x^{\star}-x_{\rm des}\|_{\infty}$, $\|x^{\star}-x_{\rm des}\|_{1}$), we solve the inverse game of finding a vector $p$ (if one exists) such that $x^{\star}\approx x_{\textrm{d}es}$.
3 KKT Conditions and Nonlinear Least-Squares Formulation
Let us assume that $f_{i}$, $g$, and $h$ are continuously differentiable with respect to $x_{i}$ and that suitable constraint qualifications hold, such as the Linear Independence Constraint Qualification (LICQ) condition at the GNE $x^{\star}(p)$ [^1] The LICQ condition is verified if the gradients of the active constraints at $x^{\star}(p)$ $\bigl\{\nabla_{x_{i}}g_{j}(x^{\star},p):j\in\mathcal{A}(x^{\star})\bigr\}\;\cup\;\bigl\{\nabla_{x_{i}}h_{k}(x^{\star},p)\bigr\}\;\cup\;\bigl\{\pm e_{k}:x_{i,k}^{\star}=\ell_{i,k}\text{ or }u_{i,k}\bigr\}$ are linearly independent, where $e_{k}$ is the $k$-th unit vector of appropriate dimension, $\mathcal{A}(x^{\star})=\{j:g_{j}(x^{\star},p)=0\}$ is the set of active inequality constraints at $x^{\star}$, and $x_{i,k}$, $\ell_{i,k}$, and $u_{i,k}$ are the $k$-th components of $x_{i}$, $\ell_{i}$, and $u_{i}$, respectively.. Then, the necessary KKT conditions of optimality for each player $i$ read as follows [32, Theorem 12.1]:
$\nabla_{x_{i}}f_{i}(x,p)+\nabla_{x_{i}}g(x,p)^{\top}\lambda_{i}+\nabla_{x_{i}}h(x,p)^{\top}\mu_{i}-v_{i}+y_{i}=0$ (3a)
$h(x,p)=0$ (3b)
$g(x,p)\leq 0,\ \ell\leq x\leq u$ (3c)
$\lambda_{i}\geq 0,\ v_{i}\geq 0,\ y_{i}\geq 0$ (3d)
$\lambda_{i,j}g_{j}(x,p)=0,\ v_{i,k}(x_{i,k}-\ell_{i,k})=0,\ y_{i,k}(u_{i,k}-x_{i,k})=0$ (3e)
$j=1,\ldots,n_{g},\ k=1,\ldots,n_{i},\ i=1,\dots,N$
where $\lambda_{i}\in\mathbb{R}^{m_{i}}$ denotes the vector of Lagrange multipliers of player $i$ associated with the shared inequality constraints $g(x,p)\leq 0$, $\mu_{i}$ the multipliers corresponding to the shared equality constraints $h(x,p)=0$, and $v_{i}$, $y_{i}$ the multipliers associated with bound constraints on $x_{i}$.
A generalized Nash equilibrium is a point $x^{\star}$ for which the KKT system (3) holds for all players simultaneously. The KKT conditions can be expressed as the following zero-finding problem:
$\begin{array}[]{rcl}0&=&\nabla_{x_{i}}f_{i}(x,p)+\nabla_{x_{i}}g(x,p)^{\top}\lambda_{i}+\nabla_{x_{i}}h(x,p)^{\top}\mu_{i}\\ 0&=&h(x,p)\\ 0&=&\sqrt{\lambda_{i,j}^{2}+g_{j}(x,p)^{2}}-\lambda_{i,j}+g_{j}(x,p)\\ 0&=&\sqrt{v_{i,k}^{2}+(\ell_{i,k}-x_{i,k})^{2}}-v_{i,k}+\ell_{i,k}-x_{i,k}\\ 0&=&\sqrt{y_{i,k}^{2}+(x_{i,k}-u_{i,k})^{2}}-y_{i,k}+x_{i,k}-u_{i,k}\\ &&k=1,\ldots,n_{i},\ i=1,\dots,N,\ j=1,\ldots,n_{g}\end{array}$ (4)
where the last three residuals encode the primal feasibility (3c), dual feasibility (3d), and complementary slackness conditions (3e) using the Fischer–Burmeister nonlinear complementarity problem (NCP) function $\phi:\mathbb{R}\to\mathbb{R}$ [18, 15] componentwise
$\phi(\alpha,\beta)\triangleq\sqrt{\alpha^{2}+\beta^{2}}-\alpha-\beta=0\quad\Longleftrightarrow\quad\alpha\beta=0,\ \alpha\geq 0,\ \beta\geq 0.$
By letting $z$ be the vector collecting $x$, $\{\lambda_{i}\}$, $\{\mu_{i}\}$, $\{v_{i}\}$, and $\{y_{i}\}$, we can express the joint KKT conditions as
$R(z,p)=0$ (5)
where $R(z,p)$ is the residual function defined in (4). In the sequel, we will denote by $\nu=\mathop{\rm col}\nolimits(\lambda,\mu,v,y)$ the stacked vector of all dual variables, so that $z=\mathop{\rm col}\nolimits(x,\nu)$ is the overall vector of primal and dual variables. Given a solution $z^{\star}(p)$ satisfying (5), we will implicitly denote by $x^{\star}(p)$ the corresponding GNE is obtained by extracting from $z^{\star}(p)$ the components corresponding to the primal variables. We will also denote by $R(z)$ the residual function when $p$ is omitted or fixed.
For conditions of existence of a GNE for a given $p$, such as nonemptiness and compactness of the set of feasible decision vectors, convexity assumptions on $f$, $g$, and linearity of $h$, and constraint qualifications, we refer the reader to [17].
3.1 Determining a GNE via nonlinear least-squares
For a given parameter $p_{0}$, we can compute the corresponding $z^{\star}(p_{0})$ by minimizing the squared residual
$z^{\star}(p_{0})\in\arg\min_{z}\frac{1}{2}\|R(z,p_{0})\|_{2}^{2}$ (6)
that serves as a merit function for the zero-finding problem (5[15, 12]. The nonlinear least-squares problem (6) can be solved via standard methods, such as Levenberg-Marquardt (LM) algorithms [28, 31], possibly exploiting sparsity in the Jacobian of $R(z,p_{0})$ with respect to $z$. If the minimum residual $R(z^{\star}(p_{0}),p_{0})$ of the nonlinear least-squares problem (6) is zero (or, more realistically, below a given positive tolerance), then $z^{\star}(p_{0})$ contains a GNE solution $x^{\star}(p_{0})$ and the associated Lagrange multipliers for each agent’s problem associated with the parameter vector $p_{0}$.
3.2 Variational GNE
Variational GNEs (vGNEs) can be obtained by simply enforcing that the Lagrange multipliers associated with the shared constraints are the same for all players, i.e., by replacing $\{\lambda_{i}\}$ with a single vector $\lambda$ and $\{\mu_{i}\}$ with a single vector $\mu$, which further reduces the number of variables in the zero-finding problem (5).
3.3 Game design
To solve the game-design problem (1), we can embed the KKT conditions (3) in the optimization of $p$, leading to the following mathematical program with equilibrium constraints (MPEC) [30, 21]:
[…]
where $z=\mathop{\rm col}\nolimits(x,\nu)$. Problem (7) can be solved via standard nonlinear programming methods with respect to both $z$ and $p$.
To avoid the complexity of dealing with the equilibrium constraints, in our approach we solve the game-design problem (1) by relaxing them and solving instead the following nonlinear programming problem:
$\begin{array}[]{rl}\min_{p,z}\quad&J([I\ 0]z,p)+\frac{\rho}{2}\|R(z,p)\|_{2}^{2}\\ \text{s.t. }&p\in P\end{array}$ (8)
where $\rho>0$ is a penalty parameter. By letting $\rho\to\infty$, the solution of (8) approaches that of (7). When $P$ is a hyper-box (with possibly infinite lower and/or upper bounds on some parameters), Problem (8) can be solved very efficiently via bound-constrained nonlinear optimization methods, such as L-BFGS-B [9].
The choice of the initial point $(z,p)$ and of the penalty $\rho$ are crucial to achieve convergence to a meaningful solution of (8); in particular, excessively small values of $\rho$ may lead to solutions that are far from satisfying the KKT conditions, while very large values of $\rho$ may cause the numerical solver to ignore the game-design objective $J(x,p)$. In practice, we can solve (8) with a moderate value of $\rho$ to get a value $(\bar{z}^{\star},\bar{p}^{\star})$ and then refine the GNE by solving the nonlinear least-squares problem (6) with hot-start at $(\bar{z}^{\star},\bar{p}^{\star})$.
3.4 Non-smooth regularization
When designing a parameter vector $p$, or finding an equilibrium among many existing ones for a given $p_{0}$, we can add the regularization term
$J_{\rm reg}(x,p)=\alpha_{1}\|x\|_{1}+\alpha_{2}\|x\|_{2}^{2},\quad\alpha_{1},\alpha_{2}\geq 0$
in the game-design objective $J$; for example, to find a sparse GNE solution $x^{\star}$, by choosing a sufficiently large $\alpha_{1}$. When $\alpha_{1}>0$, due to the non-differentiability of the $\ell_{1}$-norm at zero, we can reformulate the problem by splitting $x$ into positive and negative parts, $x=x_{p}-x_{m}$, with $x_{p},x_{m}\geq 0$ and apply the result in [5, Corollary 1] to minimize
$\begin{array}[]{rl}\min_{p,x_{p},x_{m},\nu}\quad&J(x_{p}-x_{m},p)+\frac{\rho}{2}\|R(\mathop{\rm col}\nolimits(x_{p}-x_{m},\nu),p)\|_{2}^{2}\\ &+\alpha_{1}\mathbf{1}^{\top}(x_{p}+x_{m})+\alpha_{2}(\|x_{p}\|_{2}^{2}+\|x_{m}\|_{2}^{2})\\ \text{s.t. }&p\in P,\quad x_{p}\geq 0,\ x_{m}\geq 0\end{array}$ (9)
where $\nu$ collects all dual variables.
In the special case we are just looking for a sparse GNE with no game-design objective (i.e., $J(x,p)\equiv 0$) and $\alpha_{2}>0$, by letting $\alpha_{3}\triangleq\sqrt{2\alpha_{2}}$, we have
$J_{\rm reg}(x,p)=$ $\alpha_{1}\left(\mathbf{1}^{\top}(x_{p}+x_{m})+\frac{\alpha^{2}}{2}(\|x_{p}\|_{2}^{2}+x_{m}\|_{2}^{2})\right)$
$=$ $\frac{1}{2}\left\|\alpha_{3}x_{p}+\frac{\alpha_{1}}{\alpha_{3}}\mathbf{1}\right\|_{2}^{2}+\frac{1}{2}\left\|\alpha_{3}x_{m}+\frac{\alpha_{1}}{\alpha_{3}}\mathbf{1}\right\|_{2}^{2}+\textrm{const.}$
Then, we can equivalently solve
$\min_{p,x_{p},x_{m},\nu}$ $\frac{1}{2}\left\|\begin{bmatrix}\alpha_{3}x_{p}+\frac{\alpha_{1}}{\alpha_{3}}\mathbf{1}\\ \alpha_{3}x_{m}+\frac{\alpha_{1}}{\alpha_{3}}\mathbf{1}\\ \sqrt{\rho}\ R(\mathop{\rm col}\nolimits(x_{p}-x_{m},\nu))\end{bmatrix}\right\|_{2}^{2}$ (10)
s.t. $p\in P,\ x_{p}\geq 0,\ x_{m}\geq 0.$
When $P$ is a hyper-box (with possibly infinite lower and/or upper bounds on some parameters), or a singleton $\{p_{0}\}$ as a special case, Problem (9) can be solved as a bound-constrained nonlinear least-squares problem, such as via a trust-region reflective algorithm [11]. Lower and upper bounds $\ell\leq x\leq u$ can be explicitly included in (9) and (10) by tightening the nonnegativity constraints on $x_{p}$ and $x_{m}$, i.e., by setting
$\max\{0,\ell\}\leq x_{p}\leq\max\{0,u\},\qquad\max\{0,-u\}\leq x_{m}\leq\max\{0,-\ell\}.$
The approach can be immediately extended to the case $J(x,p)=\frac{1}{2}\|f_{J}(x,p)\|_{2}^{2}$ for some vector function $f_{J}:\mathbb{R}^{n}\times\mathbb{R}^{n_{p}}\to n_{f}$ by extending the residual vector in (10) with $f_{J}(x_{p}-x_{m},p)$.
4 Linear Quadratic Games
Consider the following special case of (1) where the agents’ cost functions are quadratic and the constraints are linear:
$\begin{array}[]{rrl}\min_{p,x^{\star}}\quad&J(x^{\star},p)\\ \text{s.t. }&x^{\star}_{i}\in\arg\min_{x_{i}}&f_{i}(x,p)=\frac{1}{2}x^{\top}Q^{i}x+(c^{i}+F^{i}p)^{\top}x\\ &\text{s.t. }&Ax\leq b+Sp\\ &&A_{\mathrm{eq}}x=b_{\mathrm{eq}}+S_{\mathrm{eq}}p\\ &&\ell_{i}\leq x\leq u_{i}\\ &&x_{-i}=x^{\star}_{-i}\\ &&i=1,\dots,N\end{array}$ (11)
where $x^{\star}$ is the generalized Nash equilibrium of the Linear-Quadratic (LQ) game with $N$ players, $(Q^{i},c^{i},F^{i})$ define the quadratic cost function for agent $i$, with $Q^{i}=(Q^{i})^{\top}\in\mathbb{R}^{n\times n}$, and the diagonal block $Q^{i}_{ii}\succeq 0$, $c^{i}\in\mathbb{R}^{n}$, $F^{i}\in\mathbb{R}^{n\times n_{p}}$, matrices $A\in\mathbb{R}^{n_{g}\times n}$, $b\in\mathbb{R}^{n_{g}}$, $S\in\mathbb{R}^{n_{g}\times n_{p}}$ define the (shared) linear inequality constraints, $A_{\mathrm{eq}}\in\mathbb{R}^{n_{h}\times n}$, $b_{\mathrm{eq}}\in\mathbb{R}^{n_{h}}$, $S_{\mathrm{eq}}\in\mathbb{R}^{n_{h}\times n_{p}}$ define the (shared) linear equality constraints.
We consider as game-design function $J(x,p)$, if any is given, the sum of convex piecewise affine functions
$J_{\rm PWA}(x,p)=\sum_{j=1}^{n_{J}}\max_{k=1,\dots,n_{j}}D_{jk}x+E_{jk}p+h_{jk}$ (12a)
or a convex quadratic function
$J_{\rm Q}(x,p)=\frac{1}{2}[x^{T}\ p^{T}]Q_{J}\begin{bmatrix}x\\ p\end{bmatrix}+c_{J}^{T}\begin{bmatrix}x\\ p\end{bmatrix}$ (12b)
where $Q_{J}=Q_{J}^{\top}\succeq 0$, $Q_{J}\neq 0$, $Q_{J}\in\mathbb{R}^{(n+n_{p})\times(n+n_{p})}$, $c_{J}\in\mathbb{R}^{n+n_{p}}$, or the sum of both, i.e.,
$J(x,p)=J_{\rm PWA}(x,p)+J_{\rm Q}(x,p).$ (12c)
Special cases are: ($i$) $J_{\rm PWA}(x,p)=J_{\rm Q}(x,p)\equiv 0$ for no game design; ($ii$) $J_{\rm PWA}(x,p)=\|x-x_{\rm des}\|_{\infty}$ (i.e., $n_{J}=1$, $n_{1}=2n$, $D_{1}=[I\ -I]^{\top}$, $E_{1}=0$, $h_{1}=\mathop{\rm col}\nolimits(-x_{\rm des},x_{\rm des})$), or $J_{\rm PWA}(x,p)=\|x-x_{\rm des}\|_{1}$ (i.e., $n_{J}=n$, $n_{j}=2$, $D_{j,1}=e_{j}^{\top}$, $D_{j,2}=-e_{j}^{\top}$, $E_{j,1}=E_{j,2}=0$, $h_{j,1}=-x_{\rm des,j}$, $h_{j,2}=x_{\rm des,j}$), or $J_{\rm Q}(x,p)=\frac{1}{2}\|x-x_{\rm des}\|_{2}^{2}$ for inverse game problems; ($iv$) the social welfare quadratic objective $J_{\rm Q}(x,p)=\sum_{i=1}^{N}f_{i}(x,p)$.
For convenience, let us embed finite lower and upper bounds on $x$ into the inequality constraints by defining
$\bar{A}=\begin{bmatrix}A\\ I_{u}\\ -I_{\ell}\end{bmatrix},\quad\bar{b}=\begin{bmatrix}b\\ u_{u}\\ -\ell_{\ell}\end{bmatrix},\quad\bar{S}=\begin{bmatrix}S\\ 0\\ 0\end{bmatrix}$
where $I_{\ell}$ and $I_{u}$ collect the rows of the identity matrix corresponding, respectively, to finite upper and lower bound, and $\ell_{\ell}$ and $u_{u}$ are the corresponding bounds. Let $m$ be the total number of resulting inequality constraints.
The KKT conditions for each player’s optimization problem are necessary and sufficient for optimality, due to the convexity of the cost functions and constraints, and can be used to characterize the generalized Nash equilibria of the game:
$Q_{i}^{i}x+c^{i}_{i}+F^{i}_{i}p+\bar{A}^{\top}_{i}\lambda_{i}+A_{\mathrm{eq},i}^{\top}\mu_{i}=0$ (13)
$\bar{A}x\leq\bar{b}+\bar{S}p$
$A_{\mathrm{eq}}x=b_{\mathrm{eq}}+S_{\mathrm{eq}}p$
$\lambda_{i}\geq 0$
$\lambda_{i,j}(\bar{A}_{j}x-\bar{b}_{j}-\bar{S}_{j}p)=0,\quad i=1,\dots,N$
where $Q^{i}_{i}$ collects the rows of $Q^{i}$ corresponding to player $i$’s variables, $\bar{A}_{i}$ collects the columns and rows of $\bar{A}$ where $x_{i}$ is involved, $A_{\mathrm{eq},i}$ is defined similarly, and $\lambda_{i}$, $\mu_{i}$, are the Lagrange multipliers associated with the inequality and equality constraints, respectively, where the variables in $x_{i}$ appear.
By stacking the KKT conditions (13) for all players, we can rewrite the generalized Nash equilibrium conditions as a single mixed-integer linear program (MILP) using standard big-M constraints to model the complementarity conditions (see, e.g., [25, Proposition 2]):
$\begin{array}[]{rl}\min_{p,x,\lambda,\mu,\delta,\sigma}\quad&\displaystyle{\frac{1}{2}[x^{T}\ p^{T}]Q_{J}\begin{bmatrix}x\\ p\end{bmatrix}+c_{J}^{T}\begin{bmatrix}x\\ p\end{bmatrix}}+\sum_{j=1}^{n_{J}}\sigma_{j}\\ \text{s.t. }&\sigma_{j}\geq D_{jk}x+E_{jk}p+h_{jk},\quad k=1,\dots,n_{j}\\ &Q^{i}_{i}x+c^{i}_{i}+F^{i}_{i}p+\bar{A}^{\top}_{i}\lambda_{i}+A_{\mathrm{eq},i}^{\top}\mu_{i}=0\\ &\bar{A}x\leq\bar{b}+\bar{S}p\\ &A_{\mathrm{eq}}x=b_{\mathrm{eq}}+S_{\mathrm{eq}}p\\ &\lambda_{i}\geq 0,\ i=1,\ldots,N\\ &\bar{A}x-\bar{b}-\bar{S}p\leq M(1-\delta)\\ &\lambda_{i}\leq M\delta,\ i=1,\ldots,N\\ &\delta\in\{0,1\}^{m}\end{array}$ (14)
where $M\in\mathbb{R}$ is a sufficiently large positive constant, $\delta$ is a vector of binary variables introduced to model the complementarity conditions, and $\sigma$ is a vector of auxiliary variables used to define an upper-bound the objective function, $\sigma\in\mathbb{R}^{n_{J}}$. Clearly, at optimality $\sum_{i=1}^{n_{J}}\sigma_{j}=J_{\rm PWA}(x,p)$, as the first $n_{j}$ inequalities in (14) become all active. Note that using the same vector $\delta$ for all players ensures that shared constraints are active/inactive for all players simultaneously.
Problem (14) can be either solved with respect to $p,x,\lambda,\mu,\delta,\sigma$ by MILP when $J_{\rm Q}(x,p)\equiv 0$, or by mixed-integer quadratic programming (MIQP) otherwise.
4.1 Extracting multiple equilibria
For the same parameter vector $p$, multiple generalized Nash equilibria may exist that correspond to different combinations $\bar{\delta}$ of active inequality constraints, as recently explicitly characterized in [24]. To extract multiple equilibria, we can iteratively solve the MILP above after adding the following “no-good“ constraint
$\sum_{i:\bar{\delta}_{i}=1}\delta_{i}-\sum_{i:\bar{\delta}_{i}=0}\delta_{i}\leq\left(\sum_{i=1}^{m}\bar{\delta}_{i}\right)-1$ (15)
which prevents the MILP from returning the same combination $\bar{\delta}$ of active constraints in all subsequent iterations.
4.2 Variational GNE
As for the nonlinear case, a vGNE can be obtained by enforcing that the Lagrange multipliers associated with the shared constraints are the same for all players, which further reduces the number of variables in (14).
4.2.1 Proximal ADMM solver
For variational LQ games with fixed parameter $p=p_{0}$ and no game objective, we also consider the proximal-ADMM algorithm in [7] as an alternative approach to compute a vGNE. To apply the method, that only handles equality constraints, we reformulate the problem as the following LQ game with $N+1$ players:
$x_{i}^{\star}\in\arg\min_{x_{i}}$ $\frac{1}{2}x^{\prime}Q^{i}x+(c^{i})^{T}x$ (16)
s.t. $\bar{A}x+s=\bar{b},\quad A_{\mathrm{eq}}x=b_{\mathrm{eq}}$
$x_{-i}=x_{-i}^{\star},s=s^{\star},\quad i=1,\ldots,N$
$s^{\star}\in\arg\min_{s}$ $\frac{1}{2}\left\|\bar{A}x^{\star}+s-\bar{b}\right\|_{2}^{2}$
s.t. $\bar{A}x^{\star}+s-\bar{b},s\geq 0$
where, without loss of generality, we have taken $p_{0}=0$. The proximal-ADMM iterations for the reformulation (16) are as follows:
$x_{i}^{k+1}$ $\in\arg\min_{x_{i}}\frac{1}{2}x_{i}^{\prime}Q^{i}_{ii}x_{i}+(c^{i}_{i}+Q^{i}_{i,-i}\bar{x}_{-i}^{k}+(\lambda^{k})^{\top}\bar{A}_{i}+(\mu^{k})^{\top}A_{\mathrm{eq},i})x_{i}+\frac{\gamma}{2}\|x_{i}-x_{i}^{k}\|_{2}^{2}$
$+\frac{\rho}{2}(\|\bar{A}_{i}x_{i}^{k}+\bar{A}_{-i}\bar{x}_{-i}^{k}+s^{k}-\bar{b}\|_{2}^{2}+\|A_{\mathrm{eq},i}x_{i}+A_{\mathrm{eq},-i}\bar{x}_{-i}^{k}-d\|_{2}^{2})$ (17a)
$s^{k+1}$ $\in\arg\min_{s\geq 0}\frac{1+\rho}{2}\left\|\bar{A}x^{k+1}+s-\bar{b}\right\|_{2}^{2}+(\lambda^{k})^{\top}s+\frac{\gamma}{2}\|s-s^{k}\|_{2}^{2}$ (17b)
$\lambda^{k+1}$ $=\lambda^{k}+\rho(\bar{A}x^{k+1}+s^{k+1}-\bar{b})$ (17c)
$\mu^{k+1}$ $=\mu^{k}+\rho(A_{\mathrm{eq}}x^{k+1}-b_{\mathrm{eq}})$ (17d)
where $\bar{x}^{k}_{-i}$ collects the components of $x_{1}^{k+1},\ldots,x_{i-1}^{k+1},x_{i+1}^{k},\ldots,x_{N}^{k}$, $\gamma>0$ is the proximal parameter, and $\rho$ is the augmented Lagrangian parameter. The iterations (17) admit a closed-form solution: the $x_{i}$-update in (17a) is an unconstrained strongly convex quadratic program (QP) that can be solved by caching the Cholesky decompositions of $Q^{i}_{ii}+\gamma I+\rho(\bar{A}_{i}^{\top}\bar{A}_{i}+A_{\mathrm{eq},i}^{\top}A_{\mathrm{eq},i})$ for each player $i$ and then, at each iteration $k$, solving two triangular linear systems; the evaluation of $s^{k+1}$ also admits the closed-form solution
$s^{k+1}=\frac{1}{1+\rho+\gamma}\max\left((1+\rho)(\bar{b}-\bar{A}x^{k+1})+\gamma s^{k}-\mu^{k},0\right)$
as, by completing the squares, (17b) is equivalent to the projection of a vector onto the nonnegative orthant.
5 Game-Theoretic Control
5.1 Game-theoretic linear quadratic regulation
We consider a non-cooperative multi-agent linear-quadratic regulator (LQR) problem where $N$ agents control a shared discrete-time linear dynamical system
$x(t+1)$ $=Ax(t)+Bu(t),\qquad u(t)=\mathop{\rm col}\nolimits(u_{1}(t),\ldots,u_{N}(t))$ (18)
$y(t)$ $=Cx(t)$
where $x(t)\in\mathbb{R}^{n_{x}}$ is the state vector at step $t=0,1,\ldots$, $y(t)$ is the output vector, and $u(t)\in\mathbb{R}^{n_{u}}$ is the input vector partitioned among the $N$ agents, with $u_{i}(t)\in\mathbb{R}^{n_{i}}$ being the input applied by agent $i$ at time $t$, and $n_{u}=\sum_{i=1}^{N}n_{i}$. Each agent applies their control input $u_{i}(t)$ according to the state-feedback law
$u_{i}(t)=-K_{i}x(t)$
where $K_{i}\in\mathbb{R}^{n_{i}\times n_{x}}$ is the LQR gain corresponding to weight matrices $Q_{i}\in\mathbb{R}^{n_{x}\times n_{x}}\succeq 0$, such as $Q_{i}=C^{\top}Q_{yi}C$ with $Q_{yi}\succeq 0$, and $R_{i}\in\mathbb{R}^{n_{i}}$, $R_{i}\succ 0$. We approximate the desired LQR gain $K_{i}$ as a function of the other agents’ gains $K_{-i}$ by running the following Riccati-based iterations
$K_{i}^{k}$ $=(R_{i}+B_{i}^{T}X_{i}^{k}B_{i})^{-1}B_{i}^{T}X_{i}^{k}(A_{i}-B_{-i}K_{-i})$ (19)
$A_{k}$ $=A_{i}-B_{-i}K_{-i}-B_{i}K_{i}^{k}$
$X_{i}^{k+1}$ $=Q_{i}+A_{k}^{T}X_{i}^{k}A_{k}+(K_{i}^{k})^{T}R_{i}K_{i}^{k}$
from $X_{i}^{0}=Q_{i}$ for $k=1,\ldots,N_{\rm LQR}$. Then, we define each agent’s cost function as the squared Frobenius norm of the deviation from its optimal gain $K_{i}^{N_{\rm LQR}}(K_{-i})$:
$f_{i}(K_{i},K_{-i})=\|K_{i}^{N_{\rm LQR}}(K_{-i})-K_{i}\|_{F}^{2}.$
Let $K=[K_{1}^{\top}\ \cdots\ K_{N}^{\top}]^{\top}$ denote the stacked feedback gain. A Nash equilibrium is a feedback gain matrix $K^{\star}$ such that
$K_{i}^{\star}\in\arg\min_{K_{i}}f_{i}(K_{i},K_{-i}^{\star})\qquad\forall i=1,\dots,N.$
The equilibrium can be computed as the solution of a Nash equilibrium problem by using the nonlinear least-squares approach (6) (without parameter $p_{0}$), using the centralized LQR solution $K_{\rm LQR}$ associated with $(A,B,\sum_{i}Q_{i},\mathrm{blkdiag}(R_{1},\dots,R_{N}))$ as initial guess.
Alternatively, the coupled discrete-time algebraic Riccati equations can be solved using Riccati-based iterations as described in [33, Section III.B] until convergence is reached within a given tolerance.
5.2 Game-theoretic model predictive control
Consider again the discrete-time linear system (18) with $N$ agents competing to make the output $y(t)$ track a given set-point $r(t)\in\mathbb{R}^{n_{y}}$. By letting $\Delta u(t)=u(t)-u(t-1)$ denote the input increment at time $t$, each agent $i$ chooses the sequence $\Delta u_{i}\triangleq\{\Delta u_{i,k}\}_{k=0}^{T-1}$ of input increments over a prediction horizon of $T$ by solving the following standard discrete-time finite-horizon linear optimal control problem [6, 4]:
[…]
In (20), $Q_{y,i}\succeq 0$, $Q_{\Delta u,i}\succ 0$, and $\epsilon_{i}\geq 0$ is a slack variable used to soften shared output constraints (with linear penalty $q_{\epsilon,i}\geq 0$). For example, if an agent is interested in controlling only a subset $I_{i}$ of the components of the output vector with unit weights, we have $Q_{y,i}=\sum_{j\in I_{i}}e_{j}e_{j}^{\top}$.
The predicted trajectories $\{x_{k},y_{k},u_{k}\}$ in (20) satisfy the dynamic equations (18) over the prediction horizon $k=0,\ldots,T$, with initial condition $x_{0}=x(t)$ and $u_{-1}=u(t-1)$. Note that each agent optimizes the sequence of input increments $\Delta u_{i}$ and the slack variable $\epsilon_{i}$, so that constraints on inputs, input increments, and slacks are local to each agent, while output constraints are shared among all agents, each one softening them with possibly different slacks $\epsilon_{i}$. Hence, the resulting game-theoretic MPC problem is the GNE problem of finding sequences $\{\Delta u_{i}^{\star}\}_{i=1}^{N}$ (and slacks $\{\epsilon_{i}^{\star}\}_{i=1}^{N}$) optimizing (20) for all agents $i=1,\ldots,N$ simultaneously. In a receding-horizon fashion, only the first inputs $u_{i}(t)=u_{i}(t-1)+\Delta u_{i,0}^{\star}$ are applied for $i=1,\ldots,N$, and the entire procedure is repeated at $t+1$.
The number of constraints in each agent’s problem (20) can be reduced by imposing the constraints only on a shorter constraint horizon of $T_{c}<T$ steps. This can be especially useful when $T$ is large but we want to limit the number of Lagrange multipliers and binary variables involved in the MILP reformulation (14) of the GNEP, in which we remove $p$, $\sigma$, and the objective function (12a), as we are only interested in computing the GNE of the noncooperative MPC problem.
If a variational equilibrium is requested, the additional equalities enforcing common Lagrange multipliers for the shared output constraints in (20) can be imposed. In Section 6.5, we will consider an example of noncooperative MPC and show that it can deviate significantly from the cooperative and centralized MPC solution in which all moves are optimized jointly by solving a quadratic programming (QP) problem with decision variables $\Delta u=\mathop{\rm col}\nolimits(\Delta u_{1},\ldots,\Delta u_{N})$, $\epsilon=\mathop{\rm col}\nolimits(\epsilon_{1},\ldots,\epsilon_{N})$ and cost function given by the sum of all agents’ costs in (20).
6 Examples
We report several numerical tests to illustrate the methods illustrated in the previous sections for computing and designing generalized Nash equilibria. All tests are run on a Macbook Pro with Apple M4 Max chip and 64 GB RAM using JAX [8] for automatic differentiation and just-in-time compilation of the involved functions.
6.1 Linear-quadratic game
Let us first consider a simple linear-quadratic game with $N=3$ agents, each controlling two decision variables $x_{i}\in\mathbb{R}^{2}$, $i=1,2,3$, with cost functions
$f_{i}(x)=\frac{1}{2}x^{\top}x+i\ \mathbf{1}^{\top}x$ (21)
where $x=\mathop{\rm col}\nolimits(x_{1},x_{2},x_{3})\in\mathbb{R}^{6}$, and shared constraints
$\begin{bmatrix}-0.4&-0.1&-2.1&1.6&-1.8&-0.8\\ 0.5&-1.2&-1.1&-0.9&0.6&2.3\\ 0&-1.1&0.5&-0.6&0.0&1.2\\ -0.7&0&-0.9&-0.2&0.3&-1\end{bmatrix}x\leq\begin{bmatrix}1\\ 1\\ 1\\ 1\end{bmatrix}$.
The GNE is computed by solving the MILP (14) with no parameter $p$, and optimal combinations extracted by iteratively adding the no-good constraint (15), leading to the three equilibrium solutions:
[…]
These correspond to the optimal active constraint combinations $(1)$, $(1,4)$, and $(1,2,4)$, respectively. By imposing the additional constraint that the Lagrange multipliers associated with the shared constraints are the same for all agents, the vGNE
[…]
is found, corresponding to the active constraint combination $(1,4)$. The CPU time for computing each equilibrium ranges between $3.6$ and $10.3$ ms for the non-variational case and is $17.7$ ms for the variational one using the HiGHS solver [26].
For comparison, we also use the nonlinear least-squares approach (6) to compute the vGNE
[…]
in $233.9$ ms (11 LM iterations) starting from the initial guess $x^{0}=0$. Note that infinitely many different GNEs can exist for the same combination of active constraints, as clearly shown in [24]. The same vGNE is found by running the proximal ADMM approach described in Section 4.2.1, where we set $\gamma=1$, $\rho=1$, $x^{0}=0$, $\lambda^{0}=0$, which takes $14.6$ ms (229 ADMM iterations).
Next, we further compare the MILP, LM, and proximal ADMM approaches on random GNE problems of increasing size. We consider linear-quadratic games with $N$ agents, each controlling $n_{i}=2$ decision variables, with cost functions $f$ as in (21) and shared constraints defined by $n_{g}=2N$ random linear inequalities with unit right-hand side. The results are shown in Figure 1(a), where, for an increasing number $N$ of agents, we report the CPU time for computing a GNE via MILP and for computing a vGNE by LM and proximal ADMM. Computing GNEs on LQ games via MILP with HiGHS is about two orders of magnitude more efficient than by LM, and even more by using Gurobi’s the state-of-the-art MILP solver; it also outperforms proximal ADMM (which takes between 258 and 1794 iterations). Figure 1(b) shows instead the CPU time taken when fixing $N=5$ agents and increasing the number of shared constraints $n_{g}$ up to 300, showing that MILP methods degrade more in performance compared to proximal ADMM due to the increasing number of binary variables in (14).
(a) CPU time vs number $N$ of agents, under $n_{g}=2N$ shared constraints.
(b) CPU time vs number of shared constraints, with $N=5$ agents.
Figure 1: CPU time for computing a GNE via MILP vs LM for increasing number of agents and constraints.
6.2 Inverse linear-quadratic game design
We consider an inverse game-design problem for a GNEP with $N=10$ agents, each optimizing a vector $x_{i}\in\mathbb{R}^{n_{i}}$ of $n_{i}=10$ variables ($x\in\mathbb{R}^{100})$, parameterized by a vector $p\in\mathbb{R}^{n_{p}}$, with $n_{p}=5$, with $\|p\|_{\infty}\leq 100$. For a given $p$, each agent $i=1,\ldots,N$ solves the convex quadratic optimization problem
$\begin{array}[]{rl}x_{i}^{\star}(p)=\arg\min_{x_{i}}&\frac{1}{2}x^{\top}Q^{i}x+(c^{i}+F^{i}p)^{\top}x\\ \text{s.t. }&Ax\leq b+Sp\\ &A_{\mathrm{eq}}x=b_{\mathrm{eq}}+S_{\mathrm{eq}}p\\ &\ell_{i}\leq x\leq u_{i}\\ &x_{-i}=x^{\star}_{-i}(p)\\ &i=1,\dots,N\end{array}$ (22)
with $b\in\mathbb{R}^{50}$, $b_{\mathrm{eq}}\in\mathbb{R}^{5}$. All matrices and vectors are randomly generated with appropriate dimensions, with $Q^{i}\succ 0$, and the box constraints are defined by $u=-\ell=10\cdot\mathbf{1}$
To formulate the inverse game problem in an ideal setting, we first compute a reference equilibrium $x_{\rm des}$ and a corresponding parameter vector $\hat{p}$ by solving (11) with zero objective $J(x^{\star},p)=0$. Then, we solve the inverse game-design problem (14) by using either $J(x^{\star},p)=\|x^{\star}-x_{\rm des}\|_{\infty}$ (MILP problem) or $J(x^{\star},p)=\frac{1}{2}\|x^{\star}-x_{\rm des}\|_{2}^{2}$ (MIQP problem) to retrieve a parameter vector $p^{\star}$ that yields an equilibrium $x^{\star}(p^{\star})\approx x_{\rm des}$. We use Gurobi in both cases as solver.
The parameter vector $p^{\star}$ is retrieved, respectively, in $0.0394$ s by MILP with $\|x^{\star}(p^{\star})-x_{\rm des}\|_{\infty}=5.4712\cdot 10^{-13}$, and $0.0660$ s by MIQP with $\|x^{\star}(p^{\star})-x_{\rm des}\|_{2}=1.2434\cdot 10^{-14}$.
6.3 Stackelberg game
Consider a problem with $N$ agents (the followers), each controlling a decision variable $x_{i}\geq 0$, $i=1,\ldots,N$, subject to the shared constraint $\sum_{i=1}^{N}x_{i}\leq C$. Each follower $i$ minimizes:
$f_{i}(x,p)=a_{i}x_{i}^{2}+\sum_{j=1}^{N}\Gamma_{ij}x_{i}x_{j}+\pi_{i}(x,p)x_{i}$
where $a_{i}>0$ is a self-cost coefficient, $\Gamma\in\mathbb{R}^{N\times N}$ is a symmetric interaction matrix, and $\pi_{i}(x,p)$ is the nonlinear price function
$\pi_{i}(x,p)=p_{1,i}+p_{2}\left(\sum_{j=1}^{N}x_{j}\right)^{2}\!\!\!\!.$
Here $p=\mathop{\rm col}\nolimits(p_{1,1},\ldots,p_{1,N},p_{2})$ is the vector of game parameters chosen by the leader to minimize the loss function
$J(x,p)=\left(\sum_{i=1}^{N}x-D\right)^{2}+\eta\sum_{i=1}^{N}(p_{1,i}-\bar{p}_{1,i})^{2}-\rho\sum_{i=1}^{N}\pi_{i}(x,p)x_{i}$
with $p_{1,i}\in[p_{1,\min},p_{1,\max}]$ and $p_{2}\in[p_{2,\min},p_{2,\max}]$. The numerical values are set as follows: $N=8$, $C=1$, $D=0.9$, $p_{1,\min}=-5$, $p_{1,\max}=0$, $p_{2,\min}=0$, $p_{2,\max}=0.2$, $\bar{p}_{1,i}=-2$, $\eta=0.1$, $\rho=0.3$, and
$\Gamma=\begin{bmatrix}0&0.2&0&0&0.1&0&0&0\\ 0.2&0&0.1&0&0&0&0&0\\ 0&0.1&0&0.15&0&0&0&0\\ 0&0&0.15&0&0.1&0&0&0\\ 0.1&0&0&0.1&0&0.2&0&0\\ 0&0&0&0&0.2&0&0.1&0\\ 0&0&0&0&0&0.1&0&0.2\\ 0&0&0&0&0&0&0.2&0\end{bmatrix}\!\!,\qquad a=\begin{bmatrix}1.0\\ 1.5\\ 0.8\\ 1.2\\ 2.0\\ 0.9\\ 1.8\\ 1.1\end{bmatrix}\!\!$.
By setting $\rho=10^{8}$, (8) is solved with 83 LM iterations in 1.057 s starting from the initial guess $p_{1}^{(0)}=\frac{p_{1,\min}+p_{1,\max}}{2}$, $p_{2}^{(0)}=\frac{p_{2,\min}+p_{2,\max}}{2}$, and $x_{i}^{(0)}=\frac{C}{N}$. The optimal parameters and related GNE are
[…]
$p_{2}^{\star}$ $0\displaystyle=199$0
[…]
where $\sum_{i=1}^{N}x_{i}^{\star}(p^{\star})\approx 1=C$. The optimal value of the leader’s loss function is $J(x^{\star}(p^{\star}),p^{\star})=0.5637$. We observed that different initializations lead to different GNEs with different values of the leader’s loss function, due to the nonconvexity of the optimization problem (8).
6.4 LQR game
We consider a linear system as in (18) with $n_{x}=10$ states and $n_{u}=10$ inputs, each one controlled by a different agent ($n_{i}=1$ for all $i=1,\ldots,N$, $N=10$ agents), with $A,B\in\mathbb{R}^{10\times 10}$ randomly generated and scaled so that $A$ is unstable with spectral radius equal to 1.1. We consider state-weighting matrices $Q_{i}=e_{i}e_{i}^{T}$ and $R_{i}\equiv 10$.
In (19), we set $N_{\textrm{L}QR}=50$, which is a long-enough horizon length used to approximate the infinite-horizon cost, therefore providing a good approximation of the solution of the algebraic Riccati equation associated with the quadruple $((A-B_{-i}K_{-i}),B_{i},Q_{i},R_{i})$ for each agent $i$.
The problem is solved as in (6) (without parameter $p_{0}$) in 6.05 s (6 LM iterations), leading to an asymptotically closed-loop matrix $A-BK^{\star}$ with spectral radius 0.7525. For comparison, the centralized (cooperative) LQR solution leads to a spectral radius of 0.3965. Solving the same problem by using the Riccati-based iterations described in [33, Section III.B] leads to the same solution in 0.11 s (7 Riccati iterations).
Figure 2 compares the closed-loop step responses of the combined output vector $y(t)=\mathbf{1}^{\top}x(t)$ from various initial conditions $x(0)$ for the noncooperative LQR gain $K^{\star}$ and the centralized LQR gain $K_{\rm LQR}$.
Figure 2: Game-theoretic vs centralized LQR
6.5 Game-theoretic MPC
We consider a linear system as in (18) with $n_{x}=6$ states, $n_{u}=3$ inputs, and $n_{y}=3$ outputs, each one controlled by a different agent ($n_{i}=1$ for all $i=1,2,3$, $N=3$ agents), with $A,B\in\mathbb{R}^{6\times 3}$ randomly generated and scaled so that $A$ is stable with spectral radius equal to 0.95. The output matrix $C\in\mathbb{R}^{3\times 6}$ is scaled so that the steady-state gain from inputs to outputs is the identity matrix. Each agent $i$ has state-weighting matrix $Q_{i}=C^{\top}Q_{y,i}C$, with $Q_{y,i}=e_{i}e_{i}^{\top}$, $Q_{y,i}\in\mathbb{R}^{3\times 3}$, and input-increment weighting $Q_{\Delta u,i}=0.5$ for $i=1,2,3$. The prediction horizon is $T=10$ and the constraint horizon is $T_{c}=3$. Input constraints are $0\leq u_{i}\leq 4$ for all $i=1,2,3$, and output constraints are $0\leq y_{i}\leq 5$ for all $i=1,2,3$. The slack penalty is $q_{\epsilon,i}=10^{3}$ for all $i=1,2,3$.
The noncooperative MPC problem is solved at each time step $t=0,\ldots,T_{\rm sim}-1$, with $T_{\rm sim}=40$, by solving the MILP reformulation (14) of the GNEP (without parameter $p$ and objective function (12c)) using the HiGHS solver [26]. The resulting closed-loop trajectories are shown in Figure 3(a). For comparison, we also show the closed-loop trajectories obtained by solving a cooperative and centralized MPC problem at each time step $t$ by jointly optimizing all agents’ moves via the QP solver osQP [35].
The CPU time for solving each GNEP ranges between $4.76$ and $87.76$ ms (HiGHS MILP), while the CPU time for solving each centralized MPC problem ranges between $0.42$ and $1.50$ ms (QP).
(a) Game-theoretic MPC
(b) Centralized MPC
Figure 3: Closed-loop trajectories for linear MPC: comparison between game-theoretic (competitive) and centralized (cooperative) formulations.
6.6 Sparse GNE computations
Consider a generalized Nash equilibrium problem (GNEP) with $N=40$ agents, each one controlling a scalar decision variable $x_{i}$. Agents are grouped into pairs (2k1,2k)(2k-1,2k), $k=1,\ldots,20$, with the agents’ objectives defined as
$f_{i}(x)=\bigl(x_{2k-1}-x_{2k}\bigr)^{2},\ k=\left\lfloor(i+1)/2\right\rfloor.$
Clearly, if for every agent-pair $k$ the corresponding variables $x_{2k-1}=x_{2k}$, we have a Nash equilibrium $x^{\star}\in\mathbb{R}^{N}$. Among the infinitely many Nash equilibria, we want to minimize the game-design objective
$J(x,p)=\sum_{i=1}^{N}(x_{i}-x_{i}^{\mathrm{ref}})^{2}+\alpha_{1}\|x\|_{1}$
where the components of the reference vector $x^{\mathrm{ref}}_{i}=\frac{\lceil(i+1)/2\rceil}{10}$, $i=1,\ldots,N$, and $\alpha_{1}>0$ promotes sparsity in the equilibrium solution. Note that in this case there are no game parameters and $p$ is omitted.
Figure 4 shows the number of nonzeros in the computed GNE solution $x^{\star}$ and the optimal cost $J(x^{\star})$ as a function of the $\ell_{1}$-regularization parameter $\alpha_{1}$, obtained by solving Problem (10) with $\rho=10^{4}$. The CPU time is 1.4575 s on the first run due to JAX compilation time, and 0.0314 s on average on the subsequent runs. As expected, as $\alpha_{1}$ increases, the number of zeros in $x^{\star}$ increases, leading to sparser Nash equilibria, at the expense of a larger optimal cost $J(x^{\star})$. Clearly, due to the equilibrium condition $x_{2k-1}^{\star}=x_{2k}^{\star}$, the number of nonzeros in $x^{\star}$ can only decrease by two at a time as $\alpha_{1}$ increases.
Figure 4: Number of nonzeros in $x^{*}$ and optimal cost J(x)J(x*) as a function of the $\ell_{1}$-regularization parameter $\alpha_{1}$.
Figure 5 shows the CPU time required to compute the sparse GNE solution as a function of the number $N$ of agents for $\alpha_{1}=2$ using L-BFGS-B. The CPU time grows roughly linearly with $N$.
Figure 5: CPU time to compute the Nash equilibrium as a function of the number $N$ of agents.
7 Conclusions
In this paper, we have proposed different simple methods for designing and computing generalized Nash equilibria in noncooperative games with local and shared constraints. We have presented a mixed-integer linear or quadratic programming reformulation for computing all generalized Nash equilibria of a linear-quadratic game, as well as a nonlinear least-squares approach for computing generalized Nash equilibria of a rather general class of nonlinear games. We have illustrated the proposed methods on several numerical examples, including linear-quadratic games, game-theoretic LQR and MPC control, inverse games, and sparsity-promoting GNE computations. We believe that the proposed methods and the associated open-source library can be a useful tool for researchers and practitioners working in the field of game-theory and game-theoretic control and its applications to engineering and economic systems.